### v4

@@ 6 Conclusion
In conclusion, this research merely targeted an effective solution to detecting a cyber-attack in SDN-based VANETs, a vital network vulnerability. Therefore, even after taking note of the limitation of the dataset, the proposed ensemble-based ML model using Random Forest (RF) in combination with XGBoost and Logistic Regression has adequate accuracy in attack detection and the ability to predict with low false positives. This demonstrates that the model has a higher accuracy rate and level than the other available models, specifically in relation to the NN. Besides the above considerations, many deep learning approaches are costly and time-consuming. The proposed method presents a faster solution to detecting an attack on autonomous vehicles. Therefore, we recommend this solution be applied within an SDN controller for VANETs to enhance network performance and monitoring in the intelligent network design. Banitalebi Dehkordi et al. (2021) Furthermore, more significant hyperparameter tuning could achieve even higher accuracy for potential total accuracy.
Future Scope
Research is an endless endeavour. The future potential of this thesis resides in its ability to identify attacks using encrypted malicious data. Finally, proceed with global detection and deep learning prospects. This concept would pioneer the development of a deployable IDS solution for SDN-based VANETs. Moreover, it is possible to implement this model in other public datasets or make an extensive public synthetic VANET database with known attacks to make the model more universal and test it in the simulator before conducting a field demonstration.
The main advantage of the ensemble learning IDS is the fact that several algorithms accustomed to detecting noticeable network abnormalities are combined. One may consider the need for an algorithm to detect time-related abnormalities in the complications of DoS assailment to VANETs; others may indicate appropriate spatial irregularities that imply location spoofing. Finally, a model requires a dynamic updating mechanism, such as self-learning intrusion detection, that can adjust to new categories of threats. This constraint may be met by incorporating real-time databases and continuous learning guidelines. One may also revise future recognised assault scenarios with more sophisticated machine-learning strategies. Bonafide improves our discovery power by adding various attributes to the set. The system’s versatility is crucial to its operation, and a flexible framework that can render decoding algorithms promptly and encompasses a hazard definition is needed. The study would be substantial for establishing cybersecurity policies as it provides a model to be used for many purposes.
@@
Shakil Ibne Ahsan led the research with guidance and research support from other authors as part of research conducted at the University of the West of England, UK. All authors contributed to the preparation of the manuscript.
@@ Acknowledgements
Thank you to the School of Computing and Creative Technologies at the University of the West of England for providing the necessary support to enrich this study. The authors would also like to thank the reviewers for their valuable feedback, which has helped strengthen the article through peer review.

### v5

@@ 8 Conclusion
We have developed a robust Intrusion Detection System (IDS) framework specifically designed for Software-Defined Networks (SDN)-based Vehicle Ad-hoc Networks (VANETs). What makes this research stand out is our innovative use of stacked ensemble learning combined with SHAP analysis, which not only enhances detection accuracy but also significantly improves interpretability. While many existing models tend to focus on either performance or explainability, our approach successfully balances both, addressing a critical need in the field. By leveraging Random Forest and CatBoost alongside Logistic Regression, our model achieved an impressive accuracy 99.68% for binary classification and overall accuracy of 98.11% for multiclass classification on the VeReMi dataset. The integration of SHAP analysis adds an essential layer of transparency, revealing how different features interact and helping to pinpoint vulnerabilities that previous models often miss.
@@
Not applicable.
@@ Acknowledgements
Not applicable.